Evaluate each expression when x=1, y=3, and z=5.
1. 3y 2.z/5x 3. (2x+3y) 4. xy+z

Answer by nabla(475)

1. 3*3=9
2. 5/5=1
3. 2*1+3*3=2+9=11
4. 1*3+5=3+5=8